We have a diatomic ideal gas with a y of 5/2. It starts with an initial pressure of 1kPa, an initial temperature of 100 K, and an initial volume of 10 m^3

a) The gas undergoes an adiabatic compression, halving its volume. What is its new pressure?

b) What was the work done?

c) What was the heat flow?

d) Now, keeping pressure constant, heat is put into the gas, doubling the volume. How much heat is added?

e) What is the change in internal energy?
Answer #1

a)P2=P1*(V1/V2)^gamma=1000*(2)^(3.5/2.5)=2639Pa

b)W=-y*PV1*((P2/P1)^((gamma-1)/gamma)=-2.5*1000*10*((2639/1000)^((1.4-1)/1.4)-1)=-7988J

c)Heat flow=0 (adiabatic process)

d)Change in internal energy=nCv*(T2-T1)=2.5nR(T2-T1)=2.5P2(V2-V1)=2.5*2639*10=65975J

Work done=P2*(V2-V1)=26390J

Heat added=W+U=92365J